VARIATIONS IN TORSION-VIBRATION ENERGY STRUCTURE OF CH_{3}OH FROM FUNDAMENTAL, OVERTONE AND COMBINATION BANDS OF THE CH_{3}-ROCKING AND CO-STRETCHING MODES

\underline{R.M. LEES}, M. MOLLABASHI, LI-HONG XU, Department of Physical Sciences, University of New Brunswick, Saint John, NB, Canada E2L 4L5; MICHAEL LOCK, Physikalisch-Chemisches Institut, Justus Liebig University, D-35392 Giessen, Germany; B.P. WINNEWISSER, Department of Physics, The Ohio State University, 174 W. 18th Avenue, Columbus, OH 43210.

Torsion-vibration energy structures deduced from the ν_{7} (A′ CH_{3} in-plane rock), ν_{8} (A′ CO stretch), ν_{11} (A′′ CH_{3} outof-plane rock), 2ν_{7}, 2ν_{8}, ν_{7} + ν_{8} and ν_{8} + ν_{11} fundamental, overtone and combination bands of CH_{3}OH are compared and contrasted to that of the ground vibrational state. The 2ν_{7} in-plane CH_{3}-rocking overtone and the ν_{8} + ν_{11} COstretch/out-of-plane-rock combination bands have only recently been identified in the high-resolution Fourier transform spectrum, and point to systematic trends in the excited-state torsional behaviour for methanol. Torsion-vibration substate origins for the 8 states have been fitted to a 5-parameter Fourier model to characterize the energy patterns. Excitation of the ν_{8} mode has little influence on the torsional structure, but the A − E torsional energy splittings are sharply reduced with excitation of ν_{7} and inverted with excitation of ν_{11}. These changes are examined from the perspective of Hougen’s torsion-vibration interaction model for states of degenerate E vibrational parentage.^{a} The K = 0 substate energy pattern for the v_{7} and v_{11} states supports the model but with some differences in detail. The values of the K-scaling parameter ρ determining the periodicity of the torsional energies appear to vary almost linearly with the number of quanta of vibrational excitation. The changes are suggestive of substantial zero-point effects on the axial moments of inertia, with implications for the structural determination of the CH_{3} methyl top.

aJ.T. Hougen, J. Mol. Spectrosc. 207, 60-65 (2001).